Circular No. 5653

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 M. Della Valle and U. Kaeufl, European Southern Observatory, report:

“Spectrograms (range 400-800 nm, resolution ∼ 1 nm) obtained on Nov. 13.35 UT with the New Technology Telescope (+ EMMI) at La Silla, confirm this object (cf. IAUC 5651) to be a nova at maximum. Preliminary analysis of the spectra reveals a very narrow Hα emission line, flanked by a blue-shifted P-Cyg absorption superimposed on a strong continuum. The expansion velocity derived from the minimum of the P-Cyg absorption is ∼ 900 km/s. Emission lines of O I [mult. 20] and Ca I [mult. 30] (possibly blended with O I [mult. 38]), and Na I D are also visible. Iron lines of Fe II [mult. 42] are still visible in absorption.”

PERIODIC COMET SWIFT-TUTTLE (1992t) G. Paubert, Institut de Radio Astronomie Millimétrique; P. Colom,

D. Bockelée-Morvan, G. Bourgois, J. Crovisier, E. Gérard, and L. Jorda, Observatoire de Paris-Meudon; and D. Despois, Observatoire de Bordeaux, report: “We observed P/Swift-Tuttle with the IRAM 30-m millimeter telescope. On Nov. 11-12, HCN was detected through its J = 1-0 line at 89 GHz. The line area summed over the 3 hyperfine components is 0.27 ± 0.02 K km s^{−1}, corresponding to a production rate on the order of 1.5 × 10^{26} s^{−1}. CH_{3}OH was detected on Nov. 11 through several transitions around 145 GHz; the most intense, at 145.103 GHz, has an area of 0.16 ± 0.025 K km s^{−1}. We evaluate the CH_{3}OH production rate as ∼ 2 × 10^{27} s^{−1}. All lines are blueshifted by 0.5-0.7 km s^{−1}. The monitoring of the OH 18-cm lines at the Nançay radio telescope shows important day-to-day variations. From the average of the Nov. 10-12 observations, the OH production rate is ∼ 2.5 × 10^{29} s^{−1}. The 18-cm lines are consistently blueshifted by 0.8 ± 0.1 km s^{−1}. The strong line blueshifts observed testify to the important anisotropic outgassing and nongravitational forces in this comet.”

Further selected total visual magnitude estimates: Nov. 7.08 UT, 5.3 (P. Collins, Boulder, CO, 10×50 binoculars); 8.71, 5.5 (K. Sárneczky, Budapest, Hungary, 20×60 binoculars); 10.09, 5.4 (C. Spratt, Victoria, BC, 7×50 binoculars); 10.76, 5.3 (H. Lüthen, Hamburg, Germany, 10×50 binoculars); 13.09, 5.3 (C. S. Morris, Whitaker Peak, CA, 10×50 binoculars; 2^{o}.0 tail); 13.98, 5.3 (D. W. E. Green, Cambridge, MA, 12×50 binoculars).

1992 November 14 Daniel W. E. Green